OpenStudy (anonymous):

how many atoms are in 165 grams of calcium

OpenStudy (anonymous):

\[165g \left( \frac{ 1\ mol }{ 40.08\ g } \right)\left( \frac{ 6.02\times10^{23}\ atoms }{1\ mol } \right)\]